### v1

# PriVi: Towards A General-Purpose Video Model For Primate Behavior In The Wild

Abstract
Non-human primates are our closest living relatives, and analyzing their behavior is central to research in cognition, evolution, and conservation. Computer vision could greatly aid this research, but existing methods often rely on human-centric pretrained models and focus on single datasets, which limits generalization. We address this limitation by shifting from a model-centric to a data-centric approach and introduce PriVi, a large-scale primate-centric video pretraining dataset. PriVi contains 424 hours of curated video, combining 174 hours from behavioral research across 11 settings with 250 hours of diverse web-sourced footage, assembled through a scalable data curation pipeline. We pretrain V-JEPA on PriVi to learn primate-specific representations and evaluate it using a lightweight frozen classifier. Across four benchmark datasets – ChimpACT, BaboonLand, PanAf500, and ChimpBehave – our approach consistently outperforms prior work, including fully finetuned baselines, and scales favorably with fewer labels. These results demonstrate that primate-centric pretraining substantially improves data efficiency and generalization, making it a promising approach for low-label applications. Code, models, and the majority of the dataset will be made available.
1 Introduction
Figure 1: Our preprocessing pipeline for PriVi. YouTube data is filtered using a learned relevance classifier, while R&O is subsampled based on sub-datset metadata. We apply zero-shot primate detection on both to generate bounding boxes and discard empty frames. PriVi consists of 424 h of unique video complete with bounding boxes and CLIP embeddings for keyframes.
Understanding the behavior of wild animals is crucial for fields like cognition, ecology, and animal conservation, with the behavior of non-human primates being of especially high interest, due to its complexity and relation to human cognition. Advances in video recording technology could revolutionize the study of primate behavior by allowing to capture and analyze vast amounts of data. Current research protocols rely mostly on manual annotation by experts. This approach limits the amount of video material that can be used to draw scientific conclusions and is subject to human bias. Computer vision methods have the potential to provide complementary tools and objective means to assess video material, opening new avenues for behavioral research.
Consequently, there has been a growing interest and impressive progress in computer vision for both primate behavior [3, 8, 9, 34, 19, 16] and animal behavior in general [44, 38, 20]. Most studies introduce specialized models optimized for specific datasets which require considerable amounts of pretraining data [3, 35, 8]. However, for widespread adoption, the community needs models that can be shared between tasks and datasets and can be adapted to specific tasks with few labeled samples.
Recent foundation models demonstrate impressive generalization capabilities across many computer vision tasks, including – to some extent – animal behavior [51, 61]. Yet, pretraining data for those models is still very human-centric [4]. While there have been recent efforts to curate large-scale diverse animal video datasets [60, 59, 38], these datasets contain few to no primates. As a result, current video foundation models likely fall behind their full potential when analyzing primate behavior across a wide range of datasets with few labels. What is needed is both a large-scale pretraining datasets and a flexible approach for utilizing the foundation model to solve specific tasks on datasets. In this paper, we work towards filling these two gaps. Our contributions are:
  • Data curation pipeline, utilizing quality estimation from CLIP embeddings and zero-shot primate detection.
  • A 424-hour large-scale diverse pretraining dataset assembled using the data curation pipeline.
  • Frozen evaluation setup for behavior recognition that scales well to small datasets and is further improved by in-domain continual pretraining.
  • Our approach outperforms prior work across four primate behavior datasets.
2 Related Work
Self-supervised video learning
Learning useful representations from unlabeled data has been a cornerstone of machine learning research for decades [53, 25, 24, 55, 29, 43]. The recent success of large language models using masked language modeling [15, 40, 10] popularized self-supervised learning (SSL) as we know it today. There exists a plethora of SSL techniques ranging from contrastive learning/deep metric learning [41, 14] over self-distillation [11] to masked reconstruction in pixel [54, 23] or latent space [4, 47, 1, 2], with many modern approached combining several of those paradigms [39, 62].
The key ingredient for today’s foundation models – large-scale models that can be adapted to a wide range of tasks – is combining SSL with massive Web-scale datasets [49, 37, 31]. These image foundation models, most notably CLIP [41] and DINOv2 [39], caused a paradigm shift from handcrafted and task-specific architectures towards models using foundation models as frozen feature extractors for many computer vision tasks [33, 45]. Recently, there has also been considerable advancement in foundation models for video understanding with several architectures proposed based on video-caption pairs like PerceptionEncoder [5] and VideoCLIP [58], masked video autoencoding [54, 18], masked reconstruction in latent space like V-JEPA [4, 2], and combined approaches like VideoPrism [61].
Web-scale datasets made foundation models possible, but work on data-centric learning [21] shows that neither model nor data scale is sufficient: Careful curation of training data considerably improves performance. Many foundation models thus employ data engines for automated data curation [39, 5, 42, 30] .
Animal behavior recognition
Recent years have seen a rapidly growing number of annotated datasets for animal behavior recognition, either for single species [44, 20] or across species [38, 13]. Primate specific datasets cover diverse settings, ranging from zoo recordings in ChimpACT [34] and ChimpBehave [19] over in-the-wild recordings using camera traps in the PanAf-family of datasets [8, 9] to drone footage in BaboonLand [16]. The size of these datasets ranges between 2 h of frame-wise annotations (ChimpACT, PanAf500) to 80 h of clip-wise annotations (PanAf20k). Behavior recognition is usually operationalized as an action classification task, e. g. classifying which actions are performed in a specific miniclip [16, 19, 8], or spatio-temporal action recognition, where actions need to be localized in videos with [34] or without [35] ground truth bounding boxes of animals given.
There has been considerable prior work for automated behavior recognition: Bain et al. [3] showed that they can discriminate two distinctive actions in wild chimpanzees using audiovisual input. Recently, several methods for more challenging behavior recognition, like distinguishing between all classes in PanAf [8] or ChimpACT [34] have been proposed [35, 7, 6]. However, each of these works focuses on a single dataset and builds on models pretrained on human-centric datasets, like Kinetics [27]. They also rely on considerable amounts of training data, which is costly to acquire.
Self-supervised learning for behavior recognition
In recent years, foundation models emerged as a promising feature extractor for behavior recognition. Previous works found that frozen evaluation of VideoPRISM models is state-of-the-art or competitive with specialized methods on ChimpACT [34], KABR [28] – a dataset of Kenyan wildlife –, as well as various datasets of lab rodents [51, 61]. Mamooler et al. [36] show that jointly finetuning a large vision-language model on multiple animal behavior datasets improves performance across all, suggesting a promising step towards unified, general-purpose models without the need for specialized architectures.
For animal behavior, unlabeled data is usually easy to acquire, while labeled data is scarce. Thus, several works explored utilizing in-domain unlabeled data and found it beneficial for animal identification [26], animal behavior analysis in the lab [57], or weakly-supervised training for behavior retrieval [48]. However, all of these methods again develop models specialized for individual datasets.
3 Methodology
Table 1: Overview of the composition of our dataset. We report estimated species and setting distributions for YT-Filtered.
Subsets PriVi
YT-Filt. R&O
Unique Hours 250 h 174 h 424 h
Genus/Family [%]
Macaques $63.1$ $14.1$ $43.0$
Chimpanzees $7.8$ $35.7$ $19.3$
Orangutans $4.1$ 0 $2.4$
Baboons $1.3$ $16.2$ $7.4$
Lemuridae <1<1 $22.7$ $9.8$
Marmosets <1<1 $11.1$ $5.1$
Others $8.1$ 0 $4.8$
Not identifiable $9.4$ 0 $5.5$
No primate $6.0$ 0 $3.5$
Setting [%]
In the wild $59.6$ $62.7$ $60.9$
Wild-like $27.8$ $22.4$ $25.6$
Indoors $4.1$ $14.6$ $8.4$
Not identifiable $8.6$ 0 $5.1$
Our approach combines domain-related self-supervised pretraining with general-purpose foundation models: Instead of training only on a specific primate dataset with limited data to learn representations for this specific task, we pretrain on a broad, large-scale primate dataset to learn representations useful for several tasks. We first describe our pretraining dataset PriVi and the data pipeline used to curate it. Then we describe our framework for primate behavior recognition using broad domain-related pretraining. It consists of the self-supervised training architecture and the frozen evaluation protocol.
3.1 PriVi: Primate-Specific Pretraining Dataset
Research and Observational Data (R&O)
To overcome the data-scarce regime, where it is best for each method to train only on their own data, we assemble a diverse pretraining dataset that spans various applications of primate behavior analysis in the wild. To do so, we pool large amounts of data from various past and ongoing behavioral ecology and animal behavior research projects. In total, we pool 721 h of raw video material from 11 source datasets. A source dataset is a collection of videos that are homogeneous in their recording location, setting, and the species visible.
Five of the source datasets contain primates in their natural habitat, two in semi-free-ranging settings, and four are of captive primates. Source datasets contain both purely observational studies as well as experimental interventions, e. g. feeding boxes, matching the diversity of animal behavior recognition tasks. Recording locations in the natural habitat were in Kirindy, Madagascar; Simenti, Senegal; Phu Khieo, Thailand; and Moyen-Bafing, Guinea. Semi-free ranging locations were Straußberg, Germany and Rocamadour, France, and captive settings were at the German Primate Center, Göttingen, Germany and the University of Zurich, Switzerland.
YouTube Data
The R&O datasets features realistic settings close to the setup of existing and future applications of primate behavior analysis. However, it still has low diversity compared to web-scale pretraining datasets. To overcome this limitation, we scrape a large dataset of YouTube videos to increase diversity. We search YouTube for playlists of videos of primates in the wild and manually select 19 playlists with long, high-quality videos. Downloading these playlists yielded a corpus of 458 h of raw video material. We find that YouTube videos are diverse but on average low quality, with many frames containing e. g. interviews with human researchers, primates in cities or otherwise undesired content. We thus apply a data filtering and curation pipeline (see Section 3.2 below).
Dataset versions: PriVi, YT-Filtered, R&O, and YT-Random
After filtering, we obtain 250 unique hours of curated YouTube videos (which we refer to as YT-Filtered) and 174 unique hours of curated research and observational data (which we refer to as R&O). Both together comprise our pretraining dataset PriVi. It consists of 720,000 three-second video snippets, which are partially overlapping, depending on the source dataset. 63 % of the snippets are from YT-Filtered and 37 % from R&O. For each video snippet, the dataset contains bounding boxes and clip embeddings for the center frame. For R&O, we provide species labels inferred from the source-dataset-specific species information and a zero-shot object detector (see Section 3.2).
We release PriVi except for one source dataset comprising 14 % of our pretraining data. This source dataset contains camera trap footage from ongoing research projects and cannot yet be shared publicly. Consult Table 1 for a breakdown of the distribution of species and settings in the dataset. For ablation purposes, we create YT-Random, a YouTube dataset of the same size as YT-Filtered, but with a random selection of videos instead of the curated set.
3.2 Data Curation Pipeline
Previous work [21, 39] has repeatedly shown that even for web-scale datasets, good data curation is crucial for model performance. On the scale of hundreds of hours of video material, manual curation is infeasible. We thus automatically filter the dataset using a relevance scoring model on latent representations and zero-shot object detectors (see Fig.˜1 for an overview).
Relevance Filtering
Unlike naturalistic videos, YouTube videos feature frequent cuts. Having cuts in our training videos is undesirable, as the scene before and after the cut might be completely different, making self-supervised modeling harder and less aligned with cut-free downstream datasets. Following [60], we detect cuts in the video by using an off-the-shelf cut detector [12]. Even after cut detection, a longer scene might contain both relevant and irrelevant parts, e. g. a primate leaving the frame after half the clip. To select only relevant video snippets, we chunk videos in 3-s snippets with a stride of 2 s, allowing for an 0.5 s overlap with the previous and next snippet each, and determine for each snippet whether it is relevant for training. We specify a list of inclusion criteria to filter only non-computer-generated videos prominently featuring primates. To allow for scalable relevance filtering, one annotator manually labeled 2,500 images based on these criteria. Automated decision of these inclusion criteria requires image-level summary information, so we train a classifier model on top of CLIP embeddings of center frames for relevance filtering. Our 2-layer MLP classifier achieves 82.8 % recall and 90.3 % precision on a held-out validation set with the threshold being optimized for higher precision than recall (ROC-AUC 95.9 %).
Detection Filtering
Another considerable problem is that many collected videos contain either mostly background or are completely empty, especially for the naturalistic recordings in R&O. This is inefficient as self-supervised training would spend most of the training compute on learning representations of background information instead of capturing fine-grained details of primate behavior. Following [60], we use a Grounding DINO [32] as zero-shot object detector and prompt it for primate bounding boxes on the chunk center frames on both R&O and YT-Filtered. Afterwards, we perform thresholding, non-maximum suppression, and discard chunks without any detections. We find zero-shot primate detection to work well enough for data preprocessing, with 82.7 mAPAll on the PanAf500 primate detection task.
Subsampling
The YT dataset is already of suitable size after the relevance classifier. In R&O, however, source datasets differ vastly in size and diversity. We thus subsample based on source dataset metadata. In the R&O dataset, we aim for a proportion of 30 % for the diverse camera trap dataset, 10 % for other less diverse wild and semi-free ranging datasets, and 3 to 6 % for datasets of captive animals with low and high scene diversity, respectively. We initially extract 3-s snippets with a chunking stride between 1 s and 3 s for R&O to ensure that we capture enough snippets per dataset. After detection filtering, we subsample based on target proportion and remaining number of chunks to achieve the desired dataset composition. Note that the final proportions deviate from the target proportions, as some source datasets did not contain enough suitable samples.
3.3 Self-Supervised Pretraining
Figure 2: Our pretraining and evaluation architecture. EMA: exponential moving average. Refer to caption trainable parameters. Refer to caption frozen parameters. A. Continual pretraining of V-JEPA [4] using PriVi. B. We perform frozen evaluation with a deep, but narrow classifier. Images from our dataset and PanAf500 [8]
We utilize the V-JEPA architecture [4], a self-supervised learning approach that does not require video–text pairs because it relies on masked reconstruction in latent space. We use its default settings.
The V-JEPA architecture consists of a context encoder $E$, a target encoder $\overline{E}$, and a predictor $P$ (see Fig.˜2). The input is a video tensor $x\in\mathbb{R}^{16\times 224\times 224}$ consisting of 16 frames and with a resolution of $224\times 224$ pixels. The input tensor is obtained by performing random cropping around the bounding boxes predicted by the zero-shot primate detector (Section 3.2). The input is tokenized by using non-overlapping $2\times 16\times 16$ patches and combined with sine-cosine positional embeddings to produce $N=1568$ tokens $X\in\mathbb{R}^{D\times N}$, with token embedding dimension $D=1024$.
During training, part of the input sequence is masked and the training objective is to predict the latent representation of masked tokens given the unmasked tokens (see Fig.˜2). Let Mask and $\text{Mask}^{C}$ be functions that split a set of tokens into the set of masked and unmasked tokens, respectively. The loss for an unlabeled data sample $X$ is
$L_{\text{JEPA}}=\left\|P\left(E\left(\text{Mask}(X)\right)\right)-\text{Mask}^{C}\left(\overline{E}(X)\right)\right\|_{1}.$ (1)
The parameters of context encoder $E$ and predictor $P$ are trained. To avoid collapse to trivial solutions, the parameters of target encoder $\overline{E}$ are not trained but computed as an exponential moving average over the past weights of the context encoder $E$ [4].
3.4 Supervised Attentive Classifier
Frozen evaluation of vision backbones is not only a means to demonstrate feature quality, it is also competitive for action recognition, particularly on small datasets [51, 61]. We aim to design an attentive classifier for primate behavior recognition on a frozen pretrained model. Existing designs are suprisingly large: V-JEPA [4] uses one self-attention block (12 M parameters) and V-JEPA2 [2] three self-attention plus one cross-attention block (49 M). We hypothesize that this is overparametrized for typical small-scale animal behavior datasets, making models prone to overfitting and instable training dynamics
To reduce parameter count while keeping the classifier sufficiently deep to prevent underfitting, we downproject from $D=1024$ to $D^{\prime}=64$ dimensions before the first self-attention block in the head (Figure 2). The small latent dimension might make it difficult for one CLS token to aggregate information about all classes (e. g. 23 classes in ChimpACT). To avoid information bottlenecks, we concatenate $C$ learned CLS tokens (one per class, initialized randomly) to the patch tokens.
In summary, given $N$ patch tokens $x_{i}$ produced by the backbone $\overline{E}$ and $C$ trainable CLS tokens $q$, predicted class probabilities $\hat{y}\in[0,1]^{C}$ are calculated as
$\tilde{x}_{i}$ $=U\cdot\text{LayerNorm}(x_{i})+b$ $i\in\{1,\dots,N\}$ (2)
$X$ $=[q_{1},\dots,q_{C},\tilde{x}_{1},\dots,\tilde{x}_{N}]$ (3)
$X^{\prime}$ $=\text{SelfAttentionBlock}^{3}(X)$ (4)
$\hat{y}_{j}$ $=\sigma(v_{j}^{\text{T}}x^{\prime}_{j}+c_{j})$ $j\in\{1,\dots,C\}$ (5)
with $\sigma$ being the softmax function for single-label classification and the sigmoid function for multi-label classification. Trainable parameters include all parameters of the self attention blocks as well as $U\in\mathbb{R}^{D^{\prime}\times D}$, $b\in\mathbb{R}^{D^{\prime}}$, $q_{j}\in\mathbb{R}^{D^{\prime}}$, $v_{j}\in\mathbb{R}^{D^{\prime}}$, and $c_{j}\in\mathbb{R}$.
4 Experiments
4.1 Datasets and Evaluation Setup
We evaluate our approach on four video datasets of primate behavior: PanAf [8], ChimpBehave [19], BaboonLand [16], and ChimpACT [34].
PanAf500 [8] consists of 125 minutes of camera trap videos from 18 field sites in tropical Africa capturing chimpanzees and gorillas. Following the established protocol [8, 46], we train and evaluate on 16-frame miniclips cropped to primates using ground truth bounding boxes. Each miniclip shows one of nine behaviors.
ChimpBehave [19] is a dataset of chimpanzees in an indoor enclosure at the Basel zoo. Videos were captured using handheld cameras and behavior recognition is evaluated on 20-frame miniclips cropped to primate bounding boxes and only featuring a single behavior. The dataset features seven behavior classes, which are a subset of the PanAf500 classes. Instead of a dedicated test set, ChimpBehave utilizes five-fold cross-validation.
BaboonLand [16] consists of 18 drone recordings of wild-living olive baboons residing in Mpala, Kenya. From 30 min of densely annotated 5.3 k resolution drone footage, Duporge et al. [16] extracted 20 h of spatio-temporal miniclips, centered on each animal. Each miniclip is annotated into one of twelve behavior classes plus an additional class for occlusions. Annotation is single-label per miniclip with majority vote over per-frame labels.
ChimpACT [34] contains 2 h of video footage of chimpanzees recorded at the Leipzig Zoo. It features frame-wise bounding boxes and multi-label behavior annotations across 23 classes. Two different evaluation protocols exist: one with [34] and one without [35] ground truth bounding boxes. We follow the evaluation protocol utilizing ground truth bounding boxes. We predict the labels for a primate $i$ at frame $j$ by sampling a 64-frame miniclip around frame $j$ and producing a crop centered at $i$’s bounding box with padding to incorporate scene context.
Evaluation Metrics
Following the established evaluation protocol [16, 8, 19], we report top-1 accuracy (Acc) and average per-class accuracy (balanced accuracy, B-Acc) for PanAf500, BaboonLand, ChimpBehave. As ChimpACT is a multi-label dataset, it uses mAP instead, following the AVA protocol [22]. In addition to mAP, which weights all classes equally, we also report weighted mean average precision (mAPw), where each class average precision (AP) is weighted by its support. Thus, we report one metric weighting all classes equally (B-Acc, mAP) and one metric weighting classes by their support (Acc, mAPw) for each dataset. BaboonLand and ChimpBehave do not contain a validation split. We therefore report ablations on ChimpACT and PanAf500 only.
4.2 Architecture and Pretraining Details
Pretraining
Following V-JEPA [4], we choose a ViT-L model with video input for $E$ and a narrow 12-layer transformer for the V-JEPA predictor $P$. $E$ has 304 M parameters, $P$ has 22 M parameters. We initialize the weights from a checkpoint pretrained on VideoMix2M, consisting of Kinetics, SomethingSomethingv2, and HowTo100M, and perform pretraining for 75 k steps with a batch size of 80. This corresponds to eight epochs on the PriVi dataset. Following Singh et al. [50], we perform warmup and train with a constant small learning rate of $1.5\text{\times}{10}^{-5}$. Training takes around 11 hours on a single node with 4 A100 GPUs.
Attentive Classifier
For the attentive classifier, we always choose three layers, resulting in 220 k parameters. When training it, we freeze the encoder $\overline{E}$. On PanAf, BaboonLand and ChimpBehave, we train for 40, 10 and 30 epochs, respectively, accounting for the different number of samples in each dataset. On ChimpACT, we train for only one epoch, as the dense annotation and miniclip sampling yields 500 k highly-redundant miniclips. We use a learning rate of $1\text{\times}{10}^{-3}$ with warmup and cosine decay.
On ChimpACT and PanAf500, we report performance at the best validation accuracy, for ChimpBehave and BabboonLand, we report performance at the end of training. To reduce noise from small datasets, we always calculate class scores on three views per test sample and train five attentive classifiers and report the average. We always sample 16 frames uniformly from the input miniclip. We use cross-entropy loss except for BaboonLand where standard protocol is EQL loss [52].
5 Results
5.1 PriVi improves performance over human-centric and in-domain pretraining
Table 2: All parts of the PriVi dataset improve behavior recognition. We ablate our dataset composition. We compare pretraining on only human-centric data (VideoMix2M, V-JEPA’s default training data) with in-domain pretraining, pretraining on uncurated YouTube videos (YT-Random) and our dataset. All results are obtained on validation sets using our frozen attentive classifier, see Fig.˜2. Cross-domain comparisons in gray.
Continual Pretrain Data Unique Hours ChimpACT PanAf500
mAP mAPw Acc B-Acc
(V-JEPA) - $32.00$ $47.88$ $84.70$ $71.95$
+ ChimpACT $1.4$ $35.86$ $51.12$ 82.48 70.82
+ PanAf500 $1.5$ 28.94 45.22 $88.50$ $78.57$
YT-Random $280.0$ $33.87$ $50.55$ $86.13$ $71.61$
YT-Filtered $250.0$ $37.88$ $53.70$ 89.17 $76.33$
R&O $174.0$ $33.01$ $49.58$ $89.35$ $73.85$
PriVi (YT-F+R&O) $424.0$ 38.75 54.32 89.65 79.95
We start our analysis by exploring how much PriVi improves action recognition for a fixed, frozen backbone and attentive readout architecture. We initialize using V-JEPA weights pretrained on VideoMix2M and then pretrain with only YT-Filtered, only R&O and both combined (i. e. PriVi). Both YT-Filtered and R&O individually already improve performance on PanAf and ChimpACT over standard V-JEPA, even though R&O shows smaller gains (Table˜2). R&O contains camera trap recordings of Chimpanzees, but no recordings of zoo-housed chimpanzees, which is mirrored in the larger performance improvements on PanAf500 compared to ChimpACT. Combining YT-Filtered and R&O to PriVi yields further gains on all metrics (Table˜2), underscoring the importance of all parts of the dataset.
Table 3: In-Domain continual pretraining further improves PriVi performance We show continual pretraining on ChimpACT and PanAf500 without labels after pretraining on PriVi. All results are obtained on validation sets using our frozen attentive classifier, see Fig.˜2. Cross-domain comparisons in gray.
Continual Pretrain Data Unique Hours ChimpACT PanAf500
mAP mAPw Acc B-Acc
PriVi $424.0$ $38.75$ $54.32$ $89.65$ $79.95$
+ ChimpACT $1.4$ 41.43 57.22 84.93 69.05
+ PanAf500 $1.5$ 32.90 48.91 90.53 87.29
Table 4: Our method outperforms prior methods. We compare V-JEPA pretrained on PriVi without and with continual in-domain pretraining (CID) on the target dataset to various baseline methods including the current state-of-the-art on each dataset. AlphaChimp solves a harder task by simultaneously predicting bounding boxes; results are our own reproduction due to incompatible evaluation protocols, see Appendix.
PanAf500 ChimpACT BaboonLand ChimpBehave
Acc B-Acc mAP mAPw Acc B-Acc Acc B-Acc
X3D [17] $80.00$ $50.35$ $27.05$ $51.60$ $64.89$ $31.41$ $89.3$ $62.8$
V-JEPA [4] $81.96$ $60.19$ $30.85$ $52.89$ $72.14$ 34.57 $94.62$ $69.06$
VideoMAEv2 [54] $92.3$ 74.8
InternVideo-L [56] $78.57$ $54.01$ $25.7$
ChimpVLM [7] $84.91$ $61.94$
AlphaChimp [35] $25.35$ $40.23$
VideoPrism-g [61] $31.5$
Ours
PriVi 86.74 62.75 39.25 58.16 75.43 $33.99$ 95.58 $71.30$
PriVi+CID 85.01 62.96 40.00 59.29 76.42 38.57 96.02 75.14
YT relevance filtering boosts performance
To measure the effect of our YouTube relevance filtering, we also compare to the baseline of randomly selected YouTube videos YT-Random. Our relevance filtering considerably improves performance (Table˜2, YT-Filtered vs. YT-Random). To our surprise, even no relevance filtering still mostly outperforms the V-JEPA baseline, suggesting that even noisy, partly primate-centric data is an improvement over VideoMix2M.
Broad PriVi pretraining outperforms in-domain continual pretraining
If pretraining on in-domain data is feasible, is it still beneficial to use PriVi or should one simply train on the in-domain data? We pretrain on videos from ChimpACT and PanAf500 without labels and compare with PriVi. Even though in-domain pretraining works, it is consistently outperformed by broad PriVi pretraining (Table˜2). We also do not observe transfer learning between ChimpACT and PanAf500, while PriVi performs well on both datasets.
In-domain unlabeled video data boosts pretraining performance further
PriVi includes only broad, domain-related videos of primates. To measure how much performance increase one could expect from including unlabeled in-domain data in the pretraining, we perform continual pretraining on the target dataset (ChimpACT, PanAf500, BaboonLand or ChimpBehave) after pretraining on PriVi. Note that this pre-training is still entirely unsupervised. We use only detected, not ground-truth bounding boxes. Continual pertraining improves performance further, especially on class-balanced metrics, with $2.68\,\%$ mAP on ChimpACT and $7.34\,\%$ balanced accuracy on PanAf500 (Table˜3).
5.2 PriVi outperforms prior methods
Next, we compare our method against current state-of-the-art methods. Pretraining on PriVi followed by in-domain continual pretraining (PRIMA+CID) surpasses the state of the art on all four datasets across both class-balanced and unbalanced metrics (Section˜5.1). Our lightweight frozen attentive classifier using 220 k parameters outperforms large specialist models like ChimpVLM with 167 M parameters and full finetuning of human-centric models like VideoMAEv2.
PriVi contributes across all datasets
When comparing V-JEPA pretrained on only human-centric data with pretraining on PriVi only, PriVi brings considerable performance improvements across all four datasets.
5.3 Ablations
Finally, we establish necessity of each component by ablating various design choices for pretraining and attentive classifier. Attentive classifiers without aggressive downprojection have orders of magnitude more parameters than our approach. To ensure a fair comparison, we explore learning rates between $1\text{\times}{10}^{-3}$ and $5\text{\times}{10}^{-5}$ and perform early stopping for each ablation.
Primate-centric cropping is beneficial
During pretraining, we crop videos to detected primate bounding boxes, because we assume that it is inefficient to spend computational resources on learning representations for background patches. Indeed, primate-centric cropping improves performance considerably on every metric compared with training on full video frames (Table˜5; w/o PrimateCrop).
All components of our attentive classifier contribute
Removing the linear downprojection (w/o downproject), reducing from three to one layers (single layer), and using one instead of $C$ class tokens (single cls token) reduces performance across all metrics (Table˜5), even though the effect of $C$ class tokens is small. Even using only one layer still achieves impressive performance with only 120 k parameters, highlighting the quality of representation produced of our pretrained model.
Our pretrained model produces useful representations
A single cross-attention layer (Cross-Att.) is commonly used for frozen evaluation to demonstrate representation quality [4, 51]. While this setup performs worse than our improved attentive classifier with more layers (except on PanAf500 balanced accuracy), it still shows good performance, demonstrating good feature quality.
Ours outperforms other deep attentive classifiers
Assran et al. [2] present a deep attentive classifier with three self-attention layers followed by one cross-attention layer (Self- & Cross-Att.). Our models outperforms this design, despite theirs having 225 times more parameters.
Table 5: All components of our architecture contribute positively. We ablate (a) pretraining and (b) frozen classification design decision and (c) compare our classifier against other common attentive classifiers. All models with pretraining on PriVi only, results reported on validation sets.
Head Params ChimpACT PanAf500
mAP mAPw Acc B-Acc
Pretraining
Ours 0.22M 38.75 54.32 89.65 79.95
w/o PrimateCrop 0.22M 32.62 47.74 85.32 72.93
Classifier
Ours 0.22M 38.75 54.32 89.65 79.95
single layer 0.12M 35.68 52.16 89.21 75.81
single cls token 0.22M 38.54 53.46 88.38 74.79
w/o downproject 37.84M 30.15 46.19 88.56 62.52
Baselines
Cross-Att. 12.62M 32.02 50.64 88.80 81.99
Self- & Cross-Att. 49.63M 34.72 51.05 89.61 74.85
6 Conclusion
We introduced both PriVi, a large-scale pretraining dataset for primate behavior recognition, along with a framework for effectively utilizing unlabeled data for primate behavior recognition. We assembled PriVi from behavioral research footage spanning various species and settings, as well as YouTube videos of primates for diversity. We propose a simple, scalable data curation pipeline based on CLIP embeddings and zero-shot detections. We pretrain a V-JEPA video model using our data and use it to verify that both data sources as well as the curation pipeline contribute positively for primate behavior recognition, surpassing the state-of-the-art on all four primate behavior datasets. Thus, domain-related pretraining yields considerable performance over using a foundation model pretrained on mostly human-centric data directly.
Our proposed animal behavior recognition framework consists of (a) self-supervised pretraining of a video model on large-scale diverse animal-centric videos, (b) optional in-domain continual pretraining on the specific target domain and (c) frozen evaluation using a narrow, but deep attentive classifier. This framework has several desirable properties: Domain-related pretraining yields good behavior recognition performance across a wide variety of datasets; continual in-domain pretraining further boosts performance for individual datasets without requiring labeled data; and our frozen evaluation setup performs well even on few labeled samples. We hope that our framework and dataset will serve as a strong starting point for the community to move from specialist models for each dataset towards unified primate behavior models, which could open new avenues for research in ethology, ecology, and conservation biology.
Limitations
Even though our datasets sets a new baseline for diversity in primate behavior datasets, it still captures only 11 different research setups. Due to the limited availability of labeled datasets, we could only evaluate on chimpanzees and baboons and all of our experiments currently utilize ground truth annotated bounding boxes.
Future work
A relevant avenue for future work would be transferring our framework to other orders and suborders of mammals and eventually unifying models across mammals, as soon as sufficient pretraining data is available. Combining our approach with language models could prove beneficial for very small datasets, as it would allow in-context learning and utilization of ethograms. Further research into efficient models for real-world deployment of primate behavior recognition is needed, e. g. by distilling large video models using datasets like PriVi.
Acknowledgments
The project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – ProjectID 454648639 – SFB 1528. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK2906 – project number 502807174. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project nib00021.